# Distributed Lloyd-Based Algorithm for Uncertainty-Aware Multi-Robot Under-Canopy Flocking

### v2

@@ VI Conclusions
In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments represented by a forest, accounting for tracking errors and measurements uncertainties. We provide sufficient conditions for the theoretical guarantees of safety, proximity maintenance and convergence, compare it with state of the art algorithms, and finally we validate our solution through experiments in the field and in dense environment such as a forest. Our approach presents some limitations; the algorithm is limited to work on a plane, i.e., the reference in the altitude for each robot is constant and equal. To ensure the convergence towards the goal region, we assume that the individual obstacles are separated by a certain minimum distance (see Proposition 1, notice that these are just sufficient conditions and not necessary). Providing to the algorithm a higher level re-planner can be beneficial to enhance performance. Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to increase the agility of the swarm.

### v3

@@ VI Conclusions
In this paper, we presented a novel distributed algorithm to swarm reliably and effectively in dense cluttered environments represented by a forest. We provide sufficient conditions for safety and proximity maintenance. We propose an adaptive strategy to compensate for tracking errors and measurement uncertainties. We compare our approach with state-of-the-art algorithms, and finally we validate our solution through simulations and experiments in the field and in dense environment such as a forest. Our approach presents some limitations; the algorithm is limited to work on a plane, i.e., the reference in the altitude for each robot is constant and equal. The convergence towards the goal depends on both the environment of operation and the uncertainties taken into account. Providing to the algorithm a higher level re-planner can be beneficial to enhance performance. Future research directions will be focused on an effective 3D extension of the Lloyd algorithm and to further increase the agility of the swarm.